a large drink cost 50 cents more than a small drink. write an expression for the total cost of 3 small drinks and two large drinks
im stuck , i wrote the problem out let ld: large drinks and sd:small drinks and i dont know what to do next

Answer by HasanSahin(52)

ld = sd + 50 and total cost is >> 3*sd + 2*ld
You can write total cost only as sd OR ld.For example let's write it for only sd:
3*sd + 2*(sd+50) = 3*sd + 2*sd + 100 = 5*sd + 100
RF.